OpenStudy (anonymous):

Find the quotient. –7 ÷ 14 over 3

OpenStudy (blurbendy):

- 7 divided by 14 / 3 We can multiply -7 by the reciprocal of 14/3 (it's the same as dividing) So we end up with - 7 * 3/14 = -21/14 = -3/2